# GUI Based Fuzzy Logic and Spatial Statistics for Unsupervised Microscopy Segmentation

### v1

@@ Abstract
Brightfield microscopy imaging of unstained live cells remains a persistent challenge due to low contrast, temporal changes in specimen phenotypes, irregular illumination, and the absence of training labels. While deep learning (DL) methods (e.g., Cellpose 3.0) achieve state-of-the-art (SOTA) performance, they require extensive labeled data and heavy computational resources, and they often fail under uneven illumination. We present the first unsupervised segmentation framework combining spatial standard deviation from local mean (SSDLM), fuzzy logic, adjusted variograms, Moran’s I, and cumulative squared shift of nodal intensity (CSSNI) to address these limitations. Unlike deep learning models, our approach requires no annotations or retraining and operates through a user-friendly GUI tailored for non-programming users. The robustness and generality were validated on three datasets, including cross-domain data. We benchmark our method against 2023–2024 SOTA models, including Cellpose 3.0 and StarDist, using a dataset of unstained myoblast images. Our method achieves a significant improvement in segmentation performance, with an IoU increase of up to 48% and statistically validated superiority ($p<0.01$, Wilcoxon signed-rank test). Expert evaluation from two biologists further supports the segmentation quality (Cohen’s $\kappa>0.75$). The proposed algorithm is lightweight, interpretable, and computationally efficient, offering a practical and effective alternative for cell segmentation in label-free microscopy. The code, the dataset and the results are available for reproducibility [^1] The article is under review process. Code or Dataset may be provided upon request..
@@
journal: CMIG
@@ 1 Introduction
Our analysis identifies that using ML/DL faces ten core challenges of existing approaches: (I) Low contrast & noise – Current SOTA models achieve IoU scores below 0.3 on unstained data (see §V), as they cannot distinguish cells from background noise without strong contrast cues. (II) Uneven illumination – Inconsistent brightness across the field of view (FOV) misleads ML/DL models. (III) Optical aberrations – Dust, air bubbles, glare, and slide imperfections introduce artifacts. (IV) Overlapping structures – Cells and tissues often overlap which complicates segmentation. (V) Annotation dependency – Microscopy data demands accurate annotation, requires expert knowledge, and is both error-prone and time-consuming. Another problem is scarcity of plenty of data. (VI) Data imbalance – Rare biological events (e.g., abnormal cells) lead to biased models that perform poorly on rare but critical features. (VII) Lack of standardization – Models trained on one dataset often fail to generalize to unseen datasets or imaging conditions due to variations in microscopes, lighting, and deviation from the sample preparation. (VIII) 2D limitations – Bright-field microscopy lacks depth information, obscuring 3D features. (IX) Training dependency – Supervised tools require extensive annotations [2], making them impractical for many live-cell applications. (X) Generalization gaps – Models trained on one dataset often fail under new imaging conditions [5], especially with varying illumination or optical setups.
@@
We present a fundamentally new framework that is found to be robust under the following assumptions: I) the environment involving optical conditions and the imaging system of bright-field microscopy is set up to achieve a nearly homogeneous image plane (defined in section §II.B); II) one of two image sets (either background or specimens) must have optical heterogeneity caused by its inherent texture. The propsed model combines:
@@
    Unsupervised segmentation with ethical advantage: Our method eliminates the need for staining or labeling, preserving cell viability and avoiding cytotoxic effects—critical for live-cell and longitudinal studies.
@@
2 Academic Discussion & Background
@@ 2.4.2 Overall Time Complexity
3 Related Work
3.1 Unsupervised Learning for Image Segmentation Based on CNN
3.2 Bayesian Statistical Model
3.3 Soft Computing Based Approaches
3.4 Unsupervised Domain Adaptation for Microscopy Images
3.5 Hyperspectral Image Segmentation
3.6 Classical and Hybrid Approaches
The traditional technique along with machine learning has tried to accomplish the unsupervised learning, competitive learning, fuzzy c-means clustering, and Gibbs random fields to improve tissue component segmentation through an iterative conditional modes (ICM) algorithm adaptation. The prominent lacks of this approach lie in limited accuracy produced, subjectivity in setting thresholds and the need for ICM algorithm adaptation [26]. Another research proposed a classical approach involving edge detection and morphological processes, called MPS-Based Image Segmentation for Bright-field Microscopy [27]. The method has been implemented in open-source software Fiji. Histogram equalisation, edge detection by the Canny edge detector and filling holes by using a maximum filter are the key steps to segment the total cell area by creating a binary image. Though it works with various cell types, it has a serious lack of handling the scenarios where the cell border is obscured. Also, the report about huge testing and comparison outcomes across diverse scenarios is absent. It requires better parameter standardisation across microscopy setups. Another work, Self-Supervised Learning (SSL 222 abbreviation replicated from publisher, ”Nature Communication”. https://www.nature.com/articles/s42003-025-08190-w) Approaches [31], advances annotation-free segmentation via optical flow-based pseudo-labeling, achieving $F_{1}$ scores of 0.77–0.88 on fluorescence images (as per the author’s claim). However, for bright-field myoblasts, it exhibits major limitations: (1) 50–60 s per image due to iterative optical flow, and (2) failure on 60% of low-contrast samples where texture features are unreliable. This hampers longitudinal studies needing speed and consistency. The error ”Either no cells found or all cells are touching the border” arises when cells are undetected or too close to image edges.
@@ 4.1 Data Acquisition
The first experiments used 10 primary real-time $1920\times 1440$ images (Dataset-1) of low-contrast myoblast cultures from our lab, featuring irregular, unstained cells with often obscured boundaries due to motion. The methodology was also tested on the LIVECell dataset (Dataset-2) [29], comprising diverse cell types and $3180$ cells. Finally, we segment the images (Dataset-3) of controllable laser traces on surface patterns and evaluate the robustness of our model across domains.
@@ 4.3 Primary Masking
In this step, the $SSDLM_{\omega}$ of the patches of all the pixels are checked. For a pixel and its patch, if its $SSDLM_{\omega}$ is found to be lower than the lower bound, the pixel is turned into black by replacing the pixel value with zero.
@@ 4.9 Model Evaluation
5 Results & Discussion
Result on Dataset-1:, the segmentation outcomes of 10 bright-field microscopy images of unstained myoblast (C2C12) cultures are presented.
@@ 5 Results & Discussion
Table 2: Cohen’s Kappa Agreement Scores
Comparison Kappa ($\kappa$)
Expert 1 vs. Expert 2 0.81
Expert 1 vs. Our Model 0.78
Expert 2 vs. Our Model 0.75
Expert vs. Cellpose 0.42
Expert vs. StarDist 0.35
Each image was segmented using the proposed method, Cellpose, and StarDist for comparison. Ground truth masks were generated via expert annotations. Table 1 summarizes the average scores across 10 images. Table 4 represents the performance scores of the model with each of the 10 primary real-time images of live cells.
The proposed method outperforms both baselines in all metrics except for slightly lower accuracy compared to Cellpose, which is due to conservative masking (low false positive rate).
To validate qualitative segmentation consistency, two domain experts (cell biologists) independently assessed the outputs from all models on the original 10 images. Segmentations were rated as acceptable or not acceptable for morphological analysis. Cohen’s Kappa ($\kappa$) was computed to assess inter-observer reliability (Table 2).
The results indicate substantial agreement between the experts and the proposed method, supporting its reliability in low-contrast, label-free microscopy conditions.
Table 3: Wilcoxon Signed-Rank Test Results
textbfComparison Metric p-value
Our Model vs. Cellpose IoU 0.004
Our Model vs. StarDist IoU 0.002
Our Model vs. Cellpose F1-score 0.002
Our Model vs. StarDist F1-score 0.005
To statistically validate the superiority of the proposed method, the Wilcoxon signed-rank test (Table 3) was performed on per-image IoU and F1 scores. The tests compared our model’s results with Cellpose and StarDist across the same 10 images.
@@
All p-values are below the 0.05 threshold, indicating that the improvements are statistically significant and not due to random chance.
For Dataset-1, visual analysis (Figure 6, 7) reveals that the proposed method handles low-contrast boundaries and uneven illumination more robustly than the baselines. StarDist and Cellpose fail in detecting shapes and show boundary leakage. Our method maintains structural fidelity in elongated or clustered cell regions by leveraging spatial statistics and fuzzy inference.
Table 5: Dataset-2: Segmentation performance metrics
Metric Mean Std
Dice 0.814431 0.078670
IoU 0.694019 0.106993
Accuracy 0.872015 0.096104
Precision 0.803413 0.112562
Recall 0.851245 0.124639
F1 Score 0.814431 0.078670
SSIM 0.502930 0.245851
Hausdorff 57.287395 34.920233
@@
For Dataset-2 and Dataset-3 Fig 8 and Fig 9 are respectively attached as the representative images. Our method preserves structural fidelity in both elongated and clustered regions by combining spatial statistics with fuzzy logic-based inference. Table 6 shows group-wise average scores and segmentation profiles. While most categories used Profile-1, four of them used Profile-2. The analysis excluded eight corrupted mask instances. The average f1 scores for 3178 instances is 0.81 $\pm$ 0.07 (Tab  5).
The application of Profile-1 to the cell lines Huh7_Phase_A10, Huh7_Phase_A11, SKOV3_Phase_H4, and SKOV3_Phase_G4 results in the following mean F1-scores per category:
Huh7_Phase_A10: 0.6689(±0.0659)0.6689\ ($\pm$ 0.0659)
Huh7_Phase_A11: 0.6415(±0.0956)0.6415\ ($\pm$ 0.0956)
SKOV3_Phase_G4: 0.4742(±0.0817)0.4742\ ($\pm$ 0.0817)
SKOV3_Phase_H4: 0.5082(±0.0729)0.5082\ ($\pm$ 0.0729)
During the analysis, the following 8 images were excluded as their masks were found to be corrupted:
A172_Phase_A7_1_01d04h00m_3.png
A172_Phase_D7_1_01d20h00m_1.png
BT474_Phase_B3_1_03d00h00m_3.png
BV2_Phase_C4_1_01d16h00m_3.png
BV2_Phase_D4_1_00d12h00m_2.png
Huh7_Phase_A11_1_00d04h00m_3.png
Huh7_Phase_A11_1_00d04h00m_4.png
SHSY5Y_Phase_D10_1_01d16h00m_4.png
@@
The combination of classical spatial statistics, fuzzy logic, and adaptive masking achieves strong unsupervised performance on a challenging modality. Both expert assessment and statistical testing confirm the model’s practical value. Hyperparameter tuning via GUI sliders enhances usability, and no pre-training is required, making it accessible for low-data environments.
@@ 6 Conclusion
This method enables unsupervised, training-free segmentation of bright-field microscopy images and is validated through standard performance metrics (Accuracy, Precision, Recall, IoU, F1-score), statistical significance testing, and expert visual assessment. While it allows calibration through an interactive interface, the segmentation process is largely automated and does not require annotated training data. This makes it well-suited for regenerative medicine applications, such as stem cell tracking, wound healing assays, and time-lapse live-cell imaging. The proposed approach supports fast, reproducible analysis in experimental and clinical microscopy workflows. Future directions encompass better precision and instance segmentation for overlapping cellular communities. Also, despite the promising results, our study has limitations pertaining to the size of the dataset and relatively low precision which would be addressed in future work.

### v2

# Semi-Unsupervised Microscopy Segmentation with Fuzzy Logic and Spatial Statistics for Cross-Domain Analysis Using a GUI

@@ Abstract
Brightfield microscopy of unstained live cells is challenging due to low contrast, dynamic morphology, uneven illumination, and lack of labels. Deep learning achieved SOTA performance on stained, high-contrast images but needs large labeled datasets, expensive hardware, and fails under uneven illumination. This study presents a low-cost, lightweight, annotation-free segmentation method by introducing one-time calibration-assisted unsupervised framework adaptable across imaging modalities and image type. The framework determines background via spatial standard deviation from the local mean. Uncertain pixels are resolved using fuzzy logic, cumulative squared shift of nodal intensity, statistical features, followed by post-segmentation denoising calibration which is saved as a profile for reuse until noise pattern or object type substantially change. The program runs as a script or graphical interface for non-programmers. The method was rigorously evaluated using IoU, F1-score, and other metrics, with statistical significance confirmed via Wilcoxon signed-rank tests. On unstained brightfield myoblast (C2C12) images, it outperformed Cellpose 3.0 and StarDist, improving IoU by up to 48% (average IoU = 0.43, F1 = 0.60). In phase-contrast microscopy, it achieved a mean IoU of 0.69 and an F1-score of 0.81 on the LIVECell dataset ($n=3178$), with substantial expert agreement ($\kappa>0.75$) confirming cross-modality robustness. Successful segmentation of laser-affected polymer surfaces further confirmed cross-domain robustness. By introducing the Homogeneous Image Plane concept, this work provides a new theoretical foundation for training-free, annotation-free segmentation. The framework operates efficiently on CPU, avoids cell staining, and is practical for live-cell imaging and biomedical applications. Code and data are available for reproducibility.*
@@
codecodefootnotetext: https://github.com/anonymusUID/01cvLiveCell01p
@@ 1 Introduction
Our analysis identifies that using ML/DL faces ten core challenges of existing approaches: (I) Low contrast & noise – Current SOTA models achieve IoU scores below 0.3 on unstained data (see §V), as they cannot distinguish cells from background noise without strong contrast cues. (II) Uneven illumination – Inconsistent brightness across the field of view (FOV) misleads ML/DL models. (III) Optical aberrations – Dust, air bubbles, glare, and slide imperfections introduce artifacts. (IV) Overlapping dynamic structures – Cells and tissues often overlap with temporal changes in shape which complicates segmentation. (V) Annotation dependency – Microscopy data demands accurate annotation, requires expert knowledge, and is both error-prone and time-consuming. Another problem is scarcity of plenty of data. (VI) Data imbalance – Rare biological events (e.g., abnormal cells) lead to biased models that perform poorly on rare but critical features. (VII) Lack of standardization – Models trained on one dataset often fail to generalize to unseen datasets or imaging conditions due to variations in microscopes, lighting, and deviation from the sample preparation. (VIII) 2D limitations – Bright-field microscopy lacks depth information, obscuring 3D features. (IX) Training dependency – Supervised tools require extensive annotations [2], making them impractical for many live-cell applications. (X) Generalization gaps – Models trained on one dataset often fail under new imaging conditions [5], especially with varying illumination or optical setups.
@@
We present a fundamentally new framework that is found to be robust under the following assumptions: I) the environment involving optical conditions and the imaging system of bright-field microscopy is set up to achieve a nearly homogeneous image plane (defined in section 2); II) one of two image sets (either background or specimens) must have optical heterogeneity caused by its inherent texture. The propsed model combines:
@@
    Unsupervised segmentation for live cells: Our method eliminates the need for staining or labeling, preserving cell viability and avoiding cytotoxic effects—critical for live-cell and longitudinal studies.
@@
2 Related Work
2.1 Unsupervised Learning for Image Segmentation Based on CNN
2.2 Bayesian Statistical Model
2.3 Soft Computing Based Approaches
2.4 Unsupervised Domain Adaptation for Microscopy Images
2.5 Hyperspectral Image Segmentation
2.6 Classical and Hybrid Approaches
The traditional technique along with machine learning has tried to accomplish the unsupervised learning, competitive learning, fuzzy c-means clustering, and Gibbs random fields to improve tissue component segmentation through an iterative conditional modes (ICM) algorithm adaptation. The prominent lacks of this approach lie in limited accuracy produced, subjectivity in setting thresholds and the need for ICM algorithm adaptation [26]. Another research proposed a classical approach involving edge detection and morphological processes, called MPS-Based Image Segmentation for Bright-field Microscopy [27]. The method has been implemented in open-source software Fiji. Histogram equalisation, edge detection by the Canny edge detector and filling holes by using a maximum filter are the key steps to segment the total cell area by creating a binary image. Though it works with various cell types, it has a serious lack of handling the scenarios where the cell border is obscured. Also, the report about huge testing and comparison outcomes across diverse scenarios is absent. It requires better parameter standardisation across microscopy setups. Another work, Self-Supervised Learning (SSL 111abbreviation replicated from publisher, ”Nature Communication”. https://www.nature.com/articles/s42003-025-08190-w) Approaches [31], advances annotation-free segmentation via optical flow-based pseudo-labeling, achieving $F_{1}$ scores of 0.77–0.88 on fluorescence images (as per the author’s claim). However, for bright-field myoblasts, it exhibits major limitations: (1) 50–60 s per image due to iterative optical flow, and (2) failure on 60% of low-contrast samples where texture features are unreliable. This hampers longitudinal studies needing speed and consistency. The error ”Either no cells found or all cells are touching the border” arises when cells are undetected or too close to image edges.
3 Methodological Background & Technical Discussion
@@ 4.1 Data Acquisition
The first experiments used 10 primary real-time $1920\times 1440$ images (Dataset-1222Dataset avilable on github link.) of low-contrast myoblast (C2C12 cell) cultures from our lab, featuring irregular, unstained cells with often obscured boundaries due to motion. The cells were phographed using bright-field microscope. The methodology was also tested on the publicly avaiable LIVECell dataset (Dataset-2333https://sartorius-research.github.io/LIVECell/) [29], which comprises diverse cell types and 3,180 cells, with a modality of phase-contrast microscopy. Finally, we segment the images (Dataset-3444Dataset avilable on github link.) of controllable laser traces on surface patterns and evaluate the robustness of our model across domains.
@@ 4.3 Primary Masking
In this step, the image background is considered a nearly homogeneous, monochromatic medium, and the $SSDLM_{\omega}$ of the patches of all the pixels are checked. This assumption is grounded in the lighting mechanism of Bright-Field microscopy; however, errors / uncertainties arising from several factors will be handled in subsequent steps. For a pixel and its patch, if its $SSDLM_{\omega}$ is found to be lower than the lower bound, the pixel is turned into black by replacing the pixel value with zero.
@@ 4.9 Model Evaluation
4.10 Statistical Analysis
Performance differences between the proposed framework and baseline models were evaluated using non-parametric statistical testing. The Wilcoxon signed-rank test was applied to per-image Intersection over Union (IoU) and F1-scores to determine whether observed improvements were statistically significant, with a significance threshold of $p<0.05$. To assess reliability of expert evaluation, inter-rater agreement between biologists and between each model and expert annotations was quantified using Cohen’s Kappa coefficient ($\kappa$), where values above 0.75 indicate substantial agreement. All statistical analyses were conducted at the per-image level to ensure fair and consistent comparisons across datasets.
5 Results
5.1 Dataset-1 (unstained myoblast C2C12 cell images, n = 10, with bright-field modality. The dimension of each image is $1920\times 1440$.)
@@ 5 Results
The proposed method achieved an average Intersection over Union (IoU) of 0.431, accuracy of 0.871, precision of 0.531, recall of 0.726, and F1 score of 0.601 (Table 1). By comparison, Cellpose achieved IoU 0.130 and StarDist 0.087. Per-image performance is listed in Table 2.
Statistical testing confirmed significance: the Wilcoxon signed-rank test (Table 3) showed all p-values << 0.01 when comparing our model against both Cellpose and StarDist.
Expert evaluation indicated substantial agreement between biologists and the proposed model (Cohen’s $\kappa=0.781$ and $0.754$ vs experts, compared to $0.421$ for Cellpose and $0.351$ for StarDist; Table 4). Representative segmentations are shown in Figures 6 and 7.
Table 3: Wilcoxon Signed-Rank Test Results
textbfComparison Metric p-value
Our Model vs. Cellpose IoU 0.004
Our Model vs. StarDist IoU 0.002
Our Model vs. Cellpose F1-score 0.002
Our Model vs. StarDist F1-score 0.005
Table 4: Cohen’s Kappa Agreement Scores
Comparison Kappa ($\kappa$)
Expert 1 vs. Expert 2 0.81
Expert 1 vs. Our Model 0.78
Expert 2 vs. Our Model 0.75
Expert vs. Cellpose 0.42
Expert vs. StarDist 0.35
@@
5.2 Dataset-2 (LIVECell, n = 3178)
The framework achieved mean Dice score of 0.8144 $\pm$ 0.0786, IoU of 0.6940 $\pm$ 0.1069, accuracy of 0.8720 $\pm$ 0.0961, precision of 0.8034 $\pm$ 0.1125, and recall of 0.8512 $\pm$ 0.1246 (Table 6). Structural similarity (SSIM) was 0.5029 $\pm$ 0.2458, and average Hausdorff distance was 57.2873 $\pm$ 34.9202. Group-wise statistics are reported in Table 5.
Most cell categories performed with Profile-1, while four categories (Huh7_Phase_A10, Huh7_Phase_A11, SKOV3_Phase_G4, SKOV3_Phase_H4) required Profile-2, yielding mean F1 scores of 0.6689 $\pm$ 0.0659, 0.6415 $\pm$ 0.0956, 0.4742 $\pm$ 0.0817, and 0.5082 $\pm$ 0.0729 respectively. Representative results are shown in Figure 8.
During the analysis, the following eight images were excluded as their masks were found to be corrupted: A172_Phase_A7_1_01d04h00m_3.png, A172_Phase_D7_1_01d20h00m_1.png, BT474_Phase_B3_1_03d00h00m_3.png, BV2_Phase_C4_1_01d16h00m_3.png, BV2_Phase_D4_1_00d12h00m_2.png, Huh7_Phase_A11_1_00d04h00m_3.png, Huh7_Phase_A11_1_00d04h00m_4.png, SHSY5Y_Phase_D10_1_01d16h00m_4.png
Table 6: Dataset-2: Segmentation performance metrics
Dice IoU Accuracy Precision Recall F1 Score SSIM Hausdorff
Mean 0.814431 0.694019 0.872015 0.803413 0.851245 0.814431 0.502930 57.287395
Std 0.078670 0.106993 0.096104 0.112562 0.124639 0.078670 0.245851 34.920233
5.3 Dataset-3 (Controllable Laser Trace)
The framework maintained consistent segmentation quality (with F1-score 0.830.900.83-0.90) on simulated laser-trace images, as shown in Figure 9.
@@ 5.3 Dataset-3 (Controllable Laser Trace)
6 Discussion
The results demonstrate that the proposed framework consistently outperforms Cellpose and StarDist on unstained brightfield images, achieving up to 48% higher IoU on Dataset-1.
All Wilcoxon signed-rank tests comparing the proposed method with Cellpose and StarDist yielded $p$-values below 0.05, confirming that the observed improvements in IoU and F1-scores were statistically significant rather than random variation. Furthermore, Cohen’s Kappa analysis indicated substantial agreement ($\kappa>0.75$) between expert annotations and the proposed framework, reinforcing the reliability of the segmentation outputs in practical biomedical applications related to hard challenge – low-contrast microscopy.
The findings validate the central concept of the Homogeneous Image Plane. In microscopy, objects appear heterogeneous due to scattering and texture, while the background remains nearly homogeneous. By explicitly exploiting this property with spatial statistics and fuzzy inference, the framework separates object heterogeneity from background homogeneity.
Compared with deep learning approaches, our method has three key advantages:
  1. 1.
    Training-free and annotation-free — no labeled data or retraining needed.
  2. 2.
    Robustness across modalities — three profiles sufficed for 3200+ images from diverse cell lines. Calibration is required only when imaging modality or specimen type changes substantially.
  3. 3.
    Accessibility — runs efficiently on CPU hardware, usable via both GUI / scripts, and suited to laboratories without GPUs or annotation resources.
A limitation of the study is the relatively small size of Dataset-1. However, this was addressed by validation on the large LIVECell dataset ($n>3000$). Precision values were somewhat lower than Cellpose, reflecting a conservative design that prioritizes recall and structural fidelity — a trade-off beneficial in regenerative medicine, where false negatives are more critical than false positives.
Future extensions include automating calibration profile selection, expanding validation to additional modalities, and integrating explainable AI for improved interpretability.
Overall, the framework introduces a novel theoretical foundation through the Homogeneous Image Plane concept and demonstrates practical robustness across microscopy modalities, offering an alternative to data-hungry deep learning methods.
@@ 7 Conclusion
This method enables robust, unsupervised, training-free segmentation of bright-field microscopy images and is validated through standard performance metrics (Accuracy, Precision, Recall, IoU, F1-score), statistical significance testing, and expert visual assessment. While it allows calibration through an interactive interface, the segmentation process is largely automated and does not require annotated training data. This makes it well-suited for regenerative medicine applications, such as stem cell tracking, wound healing assays, and time-lapse live-cell imaging. The proposed approach supports fast, reproducible analysis in experimental and clinical microscopy workflows. Future directions encompass better precision and instance segmentation for overlapping cellular communities. Also, despite the promising results, our study has limitations pertaining to the size of the dataset and relatively low precision which would be addressed in future work.